Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability of red and ace ???

Expert replies
by smanstar » Tue Sep 25, 2012 8:50 pm
Q

Two cards are drawn from a pack of card , find the probability that 1st card is red and the second is an Ace, when the 1st card is.
a) Replaced ?
b) not Replaced?
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Tue Sep 25, 2012 9:38 pm
When replaced:
probability of the 1st card to be red = 1/2
when we replaced, we will have the same set of cards
probability for an ace = 4/52
probability of both events to happen = 1/2 * 4/52 = 1/26

when not replaced:
probability of the 1st card to be red = 26/52
but the 1st card can be red ace too..
so we will divide it into two red cases..
1st one red suit without an ace, and second red ace
the probability of 1st card to be red non-ace = 24/52
the probability of 2nd card to be ace = 4/51
the probability that the 1st card is non-ace red card and second card is ace = 24/52 * 4/51
the probability of the 1st card to be red ace = 2/52
the probability of the 2nd card to be ace = 3/51
the probability that the 1st card is ace red card and second card is ace = 2/52 * 3/51

the probability that 1st card is red and the second is an Ace = the probability that the 1st card is non-ace red card and second card is ace + the probability that the 1st card is ace red card and second card is ace
= (24/52 * 4/51) + (2/52 * 3/51)
= 1/26
Last edited by Anurag@Gurome on Wed Sep 26, 2012 2:19 am, edited 1 time in total.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by smanstar » Tue Sep 25, 2012 10:08 pm
Anurag@Gurome wrote:When replaced:
probability of the 1st card to be red = 1/4
when we replaced, we will have the same set of cards
probability for an ace = 4/52
probability of both events to happen = 1/4 * 4/52 = 1/52

when not replaced:
probability of the 1st card to be red = 13/52
but the 1st card can be red ace too..
so we will divide it into two red cases..
1st one red suit without an ace, and second red ace
the probability of 1st card to be red non-ace = 12/52
the probability of 2nd card to be ace = 4/51
the probability that the 1st card is non-ace red card and second card is ace = 12/52 * 4/51
the probability of the 1st card to be red ace = 1/52
the probability of the 2nd card to be ace = 3/51
the probability that the 1st card is ace red card and second card is ace = 1/52 * 3/51

the probability that 1st card is red and the second is an Ace = the probability that the 1st card is non-ace red card and second card is ace + the probability that the 1st card is ace red card and second card is ace
= 12/52 * 4/51 + 1/52 * 3/51
= 1/52



Sir

Can you explain how did you find the probability of red ?
Join the discussion

by anuprajan5 » Tue Sep 25, 2012 10:21 pm
Anurag@Gurome wrote:When replaced:
probability of the 1st card to be red = 1/4

when not replaced:
probability of the 1st card to be red = 13/52
Anurag,

I am a bit stumped with this one. There are 26 red and black cards each in a deck. Therefore, probability of the 1st card to be red should be 26/52 = 1/2

Regards
Anup
Join the discussion

by Anurag@Gurome » Wed Sep 26, 2012 2:14 am
Anurag,

I am a bit stumped with this one. There are 26 red and black cards each in a deck. Therefore, probability of the 1st card to be red should be 26/52 = 1/2

Regards
Anup
Its been long time i played cards. i was thinking of Red diamonds. Yeah the probability for a red is 1/2.
i will edit the ans.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion